OpenStudy (anonymous):

Express the number 98 as a sum of two prime numbers.

sam (.sam.):

37+61

OpenStudy (anonymous):

19 + 79 = 98 31 + 67 = 98 37 + 61 = 98 61 + 37 = 98 67 + 31 = 98 79 + 19 = 98